# Compression fronts from fast radio bursts

### v1

@@ 8.2 Unsaturated compression $(r<r_{\star})$
The saturated (steady-state) compression factor $C_{\star}$ may be reached only when the wave packet has propagated for a time $t>t_{\star}$. In the regular heating regime ($\tilde{\omega}_{B}<\tilde{\omega}_{B}$), the timescale $t_{\star}$ is given by Equation (74) in the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$). The corresponding timescale measured in any other frame (any $\kappa_{\rm u}$) is $t_{\star}\sim\sigma_{\rm u}a^{2}\gamma_{\rm u}^{2}T$ (assuming $a^{2}>1$). So, the relaxation timescale in our static lab frame is
$t_{\star}\sim\eta\gamma_{\rm u}a^{2}T.$ (93)
@@
$r_{\star}\approx\left(\frac{r_{e}{\cal E}\eta\gamma_{\rm u}}{m\omega^{2}}\right)^{1/3}\approx 10^{13}\,\frac{{\cal E}_{39}^{1/3}\eta_{4}^{4/9}}{\nu_{9}^{1/3}}\,{\rm cm},$ (95)
where ${\cal E}\sim LT$ is the FRB energy. This expression for $r_{\star}$ assumes $a^{2}(r_{\star})>1$, i.e. it holds as long as $r_{\star}<r_{1}$. At the transition radius $r=r_{\star}$ we find
$C(r_{\star})\approx C_{\star}(r_{\star})\approx\left[1+\frac{3}{2}\left(\sqrt{\frac{r_{1}^{2}}{r_{\star}^{2}}+1}-1\right)\right]^{2}.$ (96)
When $t_{\star}>t=r/c$, the propagating wave packet develops compression factors $C<C_{\star}$ (Figure 4). At times $\gamma_{\rm u}^{2}T\ll t\ll t_{\star}$, $C(t)$ may be approximated by
$C-1\approx(C_{\star}-1)\left(\frac{t}{t_{\star}}\right)^{\zeta}\approx\frac{9}{4}a^{2}\left(\frac{r}{\eta\gamma_{\rm u}a^{2}cT}\right)^{\zeta}.$ (97)
The numerical results in Figure 4 suggest $\zeta\approx 1/2$.
9 Stochastic heating zone
The zone of stochastic heating ($\tilde{\omega}_{B}>\tilde{\omega}$) is larger than one might expect from the ratio $\tilde{\omega}_{B}^{\rm u}/\tilde{\omega}_{\rm u}$ (Equation 88). Even when $\tilde{\omega}_{B}<\tilde{\omega}$ in the upstream plasma, the condition $\tilde{\omega}_{B}>\tilde{\omega}$ can be met inside the wave packet as a result of the fluid acceleration and compression: acceleration reduces $\tilde{\omega}$ and compression increases $\tilde{\omega}_{B}$.
@@ 9 Stochastic heating zone
These relations are exact for a steady compression front (with $C=C_{\star}$ and $\kappa=\kappa_{\star}$) and also approximately hold for unsaturated compression ($C\ll C_{\star}$ and $\kappa\ll\kappa_{\star}$), as demonstrated in Figure 5 using the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$). Equation (98) states the $\kappa$-$C$ relation viewed in any reference frame, where $\kappa_{\rm u}$ can have any value. The magnetic field is frozen in the fluid and compressed by the same factor as density:
$\frac{\tilde{B}}{\tilde{B}_{\rm u}}=\tilde{C},\qquad\frac{B}{B_{\rm u}}=C,$ (99)
where $\tilde{B}$ is measured in the local fluid rest frame, and subscript “u” indicates the upstream region ahead of the FRB. Equation (99) implies
$\tilde{\omega}_{B}\approx\frac{\kappa}{\kappa_{\rm u}}\,\tilde{\omega}_{B}^{\rm u}.$ (100)
The FRB frequency $\omega$ transforms to the local fluid frame as $\tilde{\omega}=\omega/\kappa$. Hence,
$\frac{\tilde{\omega}_{B}}{\tilde{\omega}}\approx\frac{\kappa^{2}}{\kappa_{\rm u}^{2}}\,\frac{\tilde{\omega}_{B}^{\rm u}}{\tilde{\omega}_{\rm u}}.$ (101)
@@
In the lab frame, the upstream medium (the wind) moves with $\kappa_{\rm u}\approx 2\gamma_{\rm u}\gg 1$, and Equation (101) may be re-written as
$\frac{\tilde{\omega}_{B}}{\tilde{\omega}}\approx C\,\frac{\tilde{\omega}_{B}^{\rm u}}{\tilde{\omega}_{\rm u}}=C\alpha.$ (102)
Now using Equation (88) one can see that $\tilde{\omega}_{B}>\tilde{\omega}$ when
$C>C_{\rm stoch}\approx\frac{\tilde{\omega}_{\rm u}}{\tilde{\omega}_{B}^{\rm u}}\approx 10\,\frac{r_{12}\nu_{9}}{L_{\rm w,37}}.$ (103)
The stochastic regime $\tilde{\omega}_{B}>\tilde{\omega}$ appears at $r<r_{\rm stoch}$, with the transition radius $r_{\rm stoch}$ typically inside $r_{\star}$. Its value is found by equating $C$ given by Equation (97) to $C_{\rm stoch}$. In particular, for $\zeta\approx 1/2$ we find
$r_{\rm stoch}\approx\left(\frac{4r_{e}^{2}L_{\rm w}L}{m^{2}c^{3}\omega^{4}\eta\gamma_{\rm u}T}\right)^{1/3}.$ (104)
This gives $r_{\rm stoch}\sim 10^{12}$ cm for typical parameters. When $r_{\rm stoch}>\eta^{1/3}R_{\rm LC}$, one can substitute $\gamma_{\rm u}\approx\eta^{1/3}$ and then Equation (104) implies $r_{\rm stoch}\propto L^{1/3}$. If one keeps $\zeta$ as a parameter that may differ from 1/2, one still finds $r_{\rm stoch}\propto L^{1/3}$:
$r_{\rm stoch}\approx 7\times 10^{11}\,H\,L_{42}^{1/3}\,{\rm cm},$ (105)
$H\approx\left(\frac{L_{\rm w,37}^{1/2}}{T_{\rm ms}^{\zeta}\eta_{4}^{4\zeta/3}\nu_{9}^{3-2\zeta}}\right)^{1/(3-3\zeta)}.$ (106)
The expression for $H$ omits a $\zeta$-dependent prefactor, which deviates from unity when $\zeta$ deviates from 1/2.
9.1 Heating and radiative cooling in the wave packet
@@
At radii $r<r_{\rm stoch}$, the FRB stochastically heats plasma particles to $\tilde{\gamma}\gg a$, far exceeding regular heating. In this regime, $\tilde{\gamma}$ monotonically grows in the fluid element as it crosses the wave packet (Figure 2). The growth is well described by Equation (18) with the numerical coefficient $\chi\approx 2/3$. It was derived in the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$), and a similar derivation can be repeated in any frame with $\kappa_{\rm u}\neq 1$ (the only change occurs in the expression for the gyro-frequency $\tilde{\omega}_{B}=\tilde{\omega}_{B}^{\rm u}\kappa/\kappa_{\rm u}$ (Equation 100)). Thus, we find
@@
$\tilde{\gamma}^{4/3}\,\frac{d\tilde{\gamma}}{d\xi}\approx\frac{16\pi}{3}\frac{r_{e}U\kappa^{2/3}}{m\omega^{4/3}}\left(\frac{\tilde{\omega}_{B}^{\rm u}}{\kappa_{\rm u}}\right)^{1/3},$ (107)
@@
$\tilde{\gamma}_{\rm stoch}\approx\left[\frac{3r_{e}{\cal E}}{mc\omega r^{2}}\right]^{3/7}\!\!\left(C\alpha\right)^{1/7}\approx 10^{5}\frac{{\cal E}_{39}^{3/7}C^{1/7}L_{\rm w,37}^{1/14}}{\nu_{9}^{4/7}r_{11}},$ (108)
[…]
Equation (108) holds as long as radiative losses of the plasma are negligible. The losses $\dot{\gamma}_{e}$ are given by Equation (36) (with the coefficient 2/3 replaced by 4/3 for stochastically heated particles, see B22). Losses $\dot{\gamma}_{e}$ offset the heating rate (Equation 10) when $\tilde{\gamma}$ approaches $\tilde{\gamma}_{\rm rad}=(c/r_{e})^{3/10}\tilde{\omega}_{B}^{1/10}\tilde{\omega}^{-2/5}$. Using $\tilde{\omega}=\omega/\kappa=\tilde{\omega}_{\rm u}\kappa_{\rm u}/\kappa$ and $\tilde{\omega}_{B}=\tilde{\omega}_{B}^{\rm u}\kappa/\kappa_{\rm u}$, we find
$\tilde{\gamma}_{\rm rad}\approx\left(\frac{c}{r_{e}\omega}\right)^{0.3}\!\alpha^{0.1}\,\frac{\kappa^{0.5}}{\kappa_{\rm u}^{0.2}}\approx 10^{4}\,\frac{C^{1/4}\kappa_{\rm u}^{0.3}L_{\rm w,37}^{1/20}}{\nu_{9}^{0.4}r_{11}^{0.1}}.$ (109)
@@
9.2 Plasma compression and trapping by the FRB
Once the radio wave begins to propagate through the plasma, it immediately induces significant compression $C$ (section 6.1), which will be denoted below as $C_{0}$. When propagation occurs in the zone of stochastic heating $r<r_{\rm stoch}$, the immediate compression $C_{0}$ turns out so large that further growth of $C>C_{0}$ on timescales $t\gg t_{\rm cross}$ (front relaxation toward the steady state) is irrelevant, because $t_{\rm cross}$ becomes too long. In general, large $C$ corresponds to large $\gamma_{\rm D}\approx C^{1/2}\gamma_{\rm u}$ and implies a long timescale for the plasma to cross the wave packet $t_{\rm cross}\sim\gamma_{\rm D}^{2}T$ (Equation 69). When $t_{\rm cross}$ reaches $t=r/c$, the plasma becomes trapped in the packet, surfing the radio wave. The ratio $t/t_{\rm cross}$ may be written as
$\frac{t_{\rm cross}}{t}\sim\frac{cT\gamma_{\rm D}^{2}}{r}=\frac{\gamma_{\rm D}^{2}}{\gamma_{\rm trap}^{2}}=\frac{C}{C_{\rm trap}},$ (112)
@@ 9.2 Plasma compression and trapping by the FRB
$\gamma_{\rm trap}\approx\left(\frac{r}{cT}\right)^{1/2},\qquad C_{\rm trap}\approx\frac{r}{cT\gamma_{\rm u}^{2}}.$ (113)
Let us now evaluate the compression factor $C_{0}$ that immediately develops in the FRB-wind interaction, and show that it easily reaches $C_{\rm trap}$. The calculation was set up in section 6.1 using the frame where the upstream plasma is at rest. Here, we will denote this frame ${\cal K}^{\prime}$ to distinguish it from our lab frame; quantities measured in frame ${\cal K}^{\prime}$ will be denoted with a prime. We are interested in the strong compression regime $\tilde{C}_{0}\gg 1$. Then, Equation (66) gives
$\tilde{C}_{0}\approx\kappa^{\prime}\approx\frac{3\tilde{\gamma}}{2\sigma_{\rm u}}+\sqrt{\frac{\epsilon_{\rm rad}^{\prime}}{\sigma_{\rm u}}},$ (114)
where we used $w\approx(3/2)\tilde{\gamma}$.
A proper compression factor $\tilde{C}$ corresponds to $C\approx\tilde{C}^{2}$ in the lab frame (Equation 98), and radiative losses in the lab frame are ${\cal E}_{\rm rad}\approx\kappa_{\rm u}{\cal E}_{\rm rad}^{\prime}$. So, one can use two general relations
$\kappa^{\prime}\approx C^{1/2},\qquad\epsilon_{\rm rad}^{\prime}\approx\frac{\epsilon_{\rm rad}}{\kappa_{\rm u}}.$ (115)
Radiative losses of a particle interacting with the radio wave packet are given by Equation (38) with an extra factor of 2 in the stochastic heating regime (B22). Thus, we find

### v2

@@ VIII.2 Unsaturated compression $(r<r_{\star})$
The saturated (steady-state) compression factor $C_{\max}^{\star}$ may be reached only when the wave packet has propagated for a time $t>t_{\star}\approx(1+\sigma_{\rm u})t_{\rm cross}^{\star}$ (section VI.1). The magnetar wind has $\sigma_{\rm u}\gg 1$ and moves with Lorentz factor $\gamma_{\rm u}\approx\kappa_{\rm u}/2\gg 1$. Then, the relaxation timescale $t_{\star}$ may be expressed as
$t_{\star}\approx\frac{3}{4}\eta\gamma_{\rm u}a_{\max}^{2}T,$ (91)
where we used Equation (61) for $t_{\rm cross}^{\star}$ and $\sigma_{\rm u}\gamma_{\rm u}=\eta$.
@@
$r_{\star}\approx\left(\frac{2r_{e}{\cal E}\eta\gamma_{\rm u}}{m\omega^{2}}\right)^{1/3}\approx 1.5\times 10^{13}\,\frac{{\cal E}_{39}^{1/3}\eta_{4}^{4/9}}{\nu_{9}^{2/3}}\,{\rm cm}.$ (93)
[…]
${\cal E}=\frac{L}{a_{\max}^{2}}\int_{0}^{T}a^{2}d\xi=\frac{3}{8}LT,$ (94)
where the numerical factor $3/8$ is found for the concrete shape of the radio wave packet given by Equation (7). The expression for $r_{\star}$ assumes $a_{\max}^{2}(r_{\star})>1$, i.e. it holds as long as $r_{\star}<r_{1}$. At the transition radius $r_{\star}$, we find
$C_{\max}^{\star}(r_{\star})\approx 1+\frac{r_{1}^{2}}{r_{\star}^{2}}.$ (95)
At radii $r<r_{\star}$, the propagation time $t<t_{\star}$ implies that the wave packet develops compression factors $C_{\max}(t)<C_{\max}^{\star}$ (Figure 6). In this regime, $C_{\max}$ may be approximated by
$C_{\max}-1\approx(C_{\max}^{\star}-1)\left(\frac{t}{t_{\star}}\right)^{\alpha}=a_{\max}^{2}\left(\frac{r}{ct_{\star}}\right)^{\alpha}.$ (96)
Numerical results in Figure 6 suggest $\alpha\approx 1/2$. The value of $\alpha$ depends on the assumed shape of the wave packet.
IX Stochastic heating zone
IX.1 Boundary of the stochastic heating zone
The transition from regular oscillations to stochastic heating is well defined on the $a$-$b$ plane (Figure 3). The transition curve $b=b_{\rm s}(a)$ is approximately described by Equation (26). Even when $b_{\rm u}\ll 1$ in the upstream plasma ahead of the FRB, the condition $b>b_{\rm s}$ can be met inside the wave packet where the fluid is accelerated and compressed. Compression increases $\tilde{\omega}_{B}$ and acceleration reduces $\tilde{\omega}$, so their ratio $b=\tilde{\omega}_{B}/\tilde{\omega}$ is increased.
@@ IX.1 Boundary of the stochastic heating zone
These relations are exact for a steady compression front (with $C=C_{\star}$ and $\kappa=\kappa_{\star}$) and also approximately hold for unsaturated compression, as demonstrated in Figure 5 using the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$). Equation (97) states the $\kappa$-$C$ relation viewed in other frames, where $\kappa_{\rm u}$ can have any value.
The magnetic field is frozen in the fluid and compressed by the same factor as density:
$\frac{\tilde{B}}{\tilde{B}_{\rm u}}=\tilde{C},\qquad\frac{B}{B_{\rm u}}=C.$ (98)
@@
This implies $\tilde{\omega}_{B}\approx(\kappa/\kappa_{\rm u})\,\tilde{\omega}_{B}^{\rm u}$ while the FRB frequency $\omega$ transforms to the local fluid frame as $\tilde{\omega}=\omega/\kappa$, so one finds $b\approx(\kappa/\kappa_{\rm u})^{2}b_{\rm u}$. In the lab frame, the upstream medium (the wind) moves with $\kappa_{\rm u}\approx 2\gamma_{\rm u}\gg 1$. Then, $(\kappa/\kappa_{\rm u})^{2}\approx C$ and
$b\approx Cb_{\rm u}.$ (99)
Large compression $C$ inside the radio wave packet can lead to $b>b_{\rm s}$, enabling stochastic heating even when $b_{\rm u}\ll 1$. Therefore, the zone of stochastic heating $r<r_{\rm stoch}$ extends far outside $r_{b}$ defined in Equation (86).
The boundary $r_{\rm stoch}$ is inside $r_{\star}$, and plasma compression by FRB at $r<r_{\star}$ occurs in the unsaturated regime: the peak $C_{\max}$ of $C(\xi)$ is reached at $\xi=\xi_{\star}<T/2$ (section III.2). The transition from regular oscillations to stochastic heating can only occur at $\xi<\xi_{\star}$ where $C(\xi)$ is rising. In this leading part of the wave packet, $C(\xi)$ tracks the steady-state $C_{\star}(\xi)$ (Figure 4), so
$C\approx 1+a^{2},\qquad b\approx(1+a^{2})b_{\rm u}.$ (100)
Stochastic heating is triggered where $b$ reaches $b_{\rm s}(a)$:
$(1+a^{2})b_{\rm u}\approx b_{\rm s}(a).$ (101)
This equation defines $a_{\rm stoch}$ and $C_{\rm stoch}\approx 1+a_{\rm stoch}^{2}$ at the transition from regular oscillations to stochastic heating. Figure 9 shows how fluid exposed to an FRB moves in the $a$-$b$ plane. At radii $r<r_{\rm stoch}$, the trajectory $b(a)$ reaches $b_{\rm s}(a)$, bringing the fluid to stochastic heating.
Figure 9: Blue curves with arrows show the evolution of parameter $b\equiv\tilde{\omega}_{B}/\tilde{\omega}$ in a fluid element exposed to an FRB. Fluid is initially in the upstream region where $a=0$ and $b=b_{\rm u}=r_{b}/r$ (Equation 86). Parameter $b$ increases inside the radio wave packet in response to the rising $a$ (Equation 100) and can eventually reach $b_{\rm s}$, triggering stochastic heating. The four blue curves correspond to radii $r/r_{b}=3$, 5, 10, 20. Reaching the stochastic heating regime at large $r$ requires high FRB luminosities $L$ that give $r_{\rm stoch}>r$ (Equation 105).
The transition point $b=b_{\rm s}$ is reached if the wave packet has $C_{\max}\geq C_{\rm stoch}$. Using $C_{\max}(r)$ given in Equation (96), one can state the condition $C_{\max}\geq C_{\rm stoch}$ as
$a_{\max}^{2}\left(\frac{4r}{3cT\eta\gamma_{\rm u}a_{\max}^{2}}\right)^{\alpha}\geq a_{\rm stoch}^{2},$ (102)
with $\alpha\approx 1/2$ suggested by our numerical models in section VI.1. Equation (102) defines the zone $r\leq r_{\rm stoch}$ where stochastic heating is triggered. Its boundary $r_{\rm stoch}$ is found using the known $b_{\rm s}(a)$ (Figure 3), the wind parameters $\eta$, $\gamma_{\rm u}(r)$, and $b_{\rm u}(r)$, and the known $a_{\max}(r)$ for a given FRB.
An approximate expression for $r_{\rm stoch}$ may be obtained using Equation (26) for $b_{\rm s}(a)$. Equation (101) then gives
$a_{\rm stoch}^{2}\approx\frac{1}{9b_{\rm u}^{2}}-1=\left(\frac{r}{3r_{b}}\right)^{2}-1.$ (103)
The boundary $r=r_{\rm stoch}$ of the stochastic heating zone (Equation 102) satisfies
$\frac{r_{1}}{r}\left(\frac{4r}{3cT\eta^{4/3}}\right)^{1/2}\approx\frac{r^{2}}{9r_{b}^{2}}-1,$ (104)
where we substituted $a_{\max}=r_{1}/r$ and $\gamma_{\rm u}\approx\eta^{1/3}$. For bright FRBs, $r_{\rm stoch}\gg 3r_{b}$ and Equation (104) gives
$r_{\rm stoch}\approx 10^{12}\frac{L_{\rm w,37}^{2/5}L_{42}^{1/5}}{\eta_{4}^{4/15}\nu_{9}^{6/5}T_{\rm ms}^{1/5}}\,{\rm cm}.$ (105)
This expression approximately holds for FRBs with $L>10^{40}\,L_{\rm w,37}^{-1/2}\eta_{4}^{4/3}\nu_{9}T_{\rm ms}$ erg s-1.
IX.2 Heating and radiative cooling in the wave packet
@@
At radii $r<r_{\rm stoch}$, the FRB stochastically heats plasma particles to $\tilde{\gamma}\gg a$, far exceeding the regular oscillation energy. In this regime, $\tilde{\gamma}$ monotonically grows in the fluid element as it crosses the wave packet (Figure 2). The growth is well described by Equation (25) with the numerical coefficient $\chi\approx 0.8$:
@@
$\tilde{\gamma}^{4/3}\,\frac{d\tilde{\gamma}}{d\xi}\approx 10\,\frac{r_{e}U\kappa^{2/3}}{m\omega^{4/3}}\left(\frac{\tilde{\omega}_{B}^{\rm u}}{\kappa_{\rm u}}\right)^{1/3},$ (106)
@@
$\tilde{\gamma}_{\rm stoch}\approx\left[\frac{2r_{e}{\cal E}}{mc\omega r^{2}}\right]^{3/7}\!\!b^{1/7}\approx 10^{5}\,\frac{{\cal E}_{39}^{3/7}L_{\rm w,37}^{1/14}\,C^{1/7}}{\nu_{9}^{4/7}r_{11}},$ (107)
where ${\cal E}=4\pi r^{2}\Sigma=\int 4\pi r^{2}Uc\,d\xi$ is the energy of the radio wave (isotropic equivalent), and we substituted $b=Cb_{\rm u}$ (Equation 99) and used Equation (86) for $b_{\rm u}$. One can see that $\tilde{\gamma}$ gained from stochastic heating exceeds the regular oscillation $\tilde{\gamma}\sim a$ by a factor of $\sim 10^{3}$.
Equation (107) holds if radiative losses of the heated plasma are negligible. Energy losses in the fluid frame $mc^{2}\tilde{\dot{\gamma}}_{e}=\Delta\tilde{\cal E}_{e}/\Delta\tilde{t}$ are given by Equation (D16), which we rewrite as
$\tilde{\dot{\gamma}}_{e}=\frac{3\sigma_{\rm T}\tilde{U}}{2mc}\,\tilde{\gamma}^{2}.$ (108)
The losses offset the heating rate (Equation 15) when $\tilde{\gamma}$ approaches
$\tilde{\gamma}_{\rm rad}\approx\left(\frac{c}{r_{e}\tilde{\omega}}\right)^{\!0.3}\!b^{0.1}\approx 10^{4}\,\frac{C^{1/4}\kappa_{\rm u}^{0.3}L_{\rm w,37}^{1/20}}{\nu_{9}^{0.4}r_{11}^{0.1}},$ (109)
where we used $b\approx Cb_{\rm u}$ and $\tilde{\omega}=\omega/\kappa=(\omega/\kappa_{\rm u}\sqrt{C})$.
@@
IX.3 Plasma trapping by the FRB
When FRB propagation through the wind occurs in the zone of stochastic heating $r<r_{\rm stoch}$, the plasma acceleration and compression turn out so large that it becomes trapped inside the radio wave packet. In general, large compression $C$ corresponds to large $\gamma_{\rm D}\approx C^{1/2}\gamma_{\rm u}$ (Equation 97) and implies a long timescale for the plasma to cross the wave packet $t_{\rm cross}\sim\gamma_{\rm D}^{2}T$ (Equation 60). If $t_{\rm cross}$ exceeds $t=r/c$, the plasma becomes trapped, surfing the radio wave. The ratio $t_{\rm cross}/t$ may be written as
$\frac{t_{\rm cross}}{t}\sim\frac{cT\gamma_{\rm D}^{2}}{r}\approx\frac{\gamma_{\rm D}^{2}}{\gamma_{\rm trap}^{2}}\approx\frac{C}{C_{\rm trap}},$ (112)
@@ IX.3 Plasma trapping by the FRB
$\gamma_{\rm trap}\approx\left(\frac{r}{cT}\right)^{1/2},\quad C_{\rm trap}\approx\frac{r}{cT\gamma_{\rm u}^{2}}\approx\frac{30\,r_{11}}{T_{\rm ms}(\gamma_{\rm u}/10)^{2}}.$ (113)
In the regular oscillation zone $r>r_{\rm stoch}$, trapping does not occur: $C_{\max}(r)$ is everywhere below $C_{\rm trap}(r)$. Note that $C_{\max}$ is a decreasing function of radius while $C_{\rm trap}$ increases with radius, so the ratio $C_{\max}/C_{\rm trap}$ is largest at the inner boundary of the regular oscillation zone $r_{\rm stoch}$. It is easy to verify that $C_{\max}(r_{\rm stoch})=C_{\rm stoch}<C_{\rm trap}(r_{\rm stoch})$, where $C_{\rm stoch}$ is given by Equation (103).
By contrast, in the stochastic heating zone $r<r_{\rm stoch}$, the FRB deposits enormous energy and momentum into the plasma, resulting in huge compression $C$. It would greatly exceed $C_{\rm trap}$ if one assumes that the plasma crosses the wave packet. [^12] A lower limit on the proper compression factor $\tilde{C}\approx C^{1/2}\approx\tilde{\kappa}/\tilde{\kappa}_{\rm u}$ may be estimated as $\tilde{\gamma}/\sigma_{\rm u}$, which gives $C\gg C_{\rm trap}$. So, the plasma cannot cross the packet at $r<r_{\rm stoch}$; it gets stuck inside the packet at some $\xi_{\rm trap}<T$.
The trapping regime holds unless the FRB itself becomes strongly damped by its interaction with the wind. Damping allows the plasma to move through the wave packet while consuming its energy until $\xi_{\rm trap}$ approaches $T$ and the entire FRB loses its energy. The parameter space where this happens is evaluated below.
X FRB energy losses
Consider an FRB wave packet overtaking the wind and accelerating it from Lorentz factor $\gamma_{\rm u}$ to some $\gamma_{\rm D}$. Acceleration by the factor $\gamma_{\rm u}/\gamma_{\rm D}\approx\kappa_{\rm u}/\kappa\gg 1$ is accompanied by compression given in Equation (97).